# TrAME: Trajectory-Anchored Multi-View Editing for Text-Guided 3D Gaussian Splatting Manipulation

### v1

@@ Abstract
Despite significant strides in the field of 3D scene editing, current methods encounter substantial challenge, particularly in preserving 3D consistency in multi-view editing process. To tackle this challenge, we propose a progressive 3D editing strategy that ensures multi-view consistency via a Trajectory-Anchored Scheme (TAS) with a dual-branch editing mechanism. Specifically, TAS facilitates a tightly coupled iterative process between 2D view editing and 3D updating, preventing error accumulation yielded from text-to-image process. Additionally, we explore the relationship between optimization-based methods and reconstruction-based methods, offering a unified perspective for selecting superior design choice, supporting the rationale behind the designed TAS. We further present a tuning-free View-Consistent Attention Control (VCAC) module that leverages cross-view semantic and geometric reference from the source branch to yield aligned views from the target branch during the editing of 2D views. To validate the effectiveness of our method, we analyze 2D examples to demonstrate the improved consistency with the VCAC module. Further extensive quantitative and qualitative results in text-guided 3D scene editing indicate that our method achieves superior editing quality compared to state-of-the-art methods. We will make the complete codebase publicly available following the conclusion of the double-blind review process.
1 Introduction
@@
3D technologies are pivotal in various fields, such as virtual reality (VR), the film and gaming industries, and 3D design [40, 11, 22, 12]. Recent advances in Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) have pushed the boundary of 3D representation, enabling versatile 3D applications [21, 37, 23, 31, 7, 6, 17, 35, 33, 3]. Leveraging the powerful generative capabilities of text-to-image diffusion models [2, 26], text-guided 3D editing now allows for intricate adjustments in shape, style, texture, and lighting, marking a significant advancement in the flexibility of 3D scene manipulation [9, 16, 11, 22, 12, 18].
@@
One key challenge in 3D editing is maintaining multi-view consistency when applied to real scenes [29, 32], which is essential for preventing visual artifacts and inconsistent appearances when observed from different view angles. Current studies of 3D editing can be divided into two categories, namely optimization-based and reconstruction-based methods. The optimization-based 3D editing methods [19, 27, 20, 41, 16] usually adopt the Score Distillation Sampling (SDS) loss [25] or its modified versions [13, 24]. They often suffer from sub-optimal editing quality such as over-saturation, over-smoothing, and a lack of diversity due to the misalignment between the randomly sampled timestep $t$ in the SDS optimization process and the corresponding timestep $t$ in the diffusion sampling process [15, 40]. The second category of methods, reconstruction-based 3D editing methods [9, 5, 32, 29, 18], employ readily available 2D editing diffusion models to edit 3D scene. However, 2D diffusion models struggle to achieve multi-view consistency because of their inherent limitation of independently processing each view.
@@ 2 Related Work
Regarding optimization-based 3D editing, significant advancements have been made, building upon the foundational SDS loss introduced by DreamFusion [25]. This has catalyzed the development of various innovative methods [19, 27, 20, 41, 16, 24] aimed at refining 3D model editing via optimizing the SDS loss. Notably, Vox-E [27] and DreamEditor [41] have harnessed explicit 3D representations, including voxels and meshes. These approaches utilize cross-attention mechanisms to facilitate precise manipulations within specific regions of 3D models. Furthermore, RePaint-NeRF [39] has advanced the application of SDS in 3D editing by integrating a semantic mask to guide and constrain modifications within the background elements. In a similar vein, ED-NeRF [24] has introduced an enhanced loss function specifically designed for 3D editing tasks. It extends the Delta Denoising Score (DDS) loss [13] into the three-dimensional space, offering a more refined approach to editing. Despite these advancements, a common challenge still persists among methods that utilize randomly sampled timesteps $t$ within the SDS framework: This approach tends to lead to deviations from the intended diffusion sampling trajectories, which can result in compromised quality of the 3D edits.
For reconstruction-based 3D editing, InstructN2N [9] alongside subsequent studies [5, 32, 29, 18], have leveraged advancements in 2D diffusion editing techniques, notably InstructP2P [2] and ControlNet [38], for enhancing scene updates through Iterative Dataset Update [9]. GaussianEditor [5] and GaussCtrl [32] introduce methods for precise control in partial 3D editing. GaussianEditor utilizes semantic tracing to segment the 3D Gaussians to be edited. GaussCtrl proposes depth-guided editing and attention-based latent alignment. However, these methods often neglect inter-view correlations, resulting in a gap in maintaining semantic and structural consistency across multiple views, ultimately leading to error accumulation. Our work bridges this gap by integrating Trajectory-Anchored Scheme to minimize error accumulation and introducing attention manipulation mechanisms to enforce structural and semantic consistency across views in 3D editing through our VCAC module.
3 Preliminaries
@@
3.1 Score Distillation Sampling
@@
3.2 Denoising Diffusion Consistent Model
@@
${\bm{z}}_{t-1}=\sqrt{\alpha_{t-1}}\underbrace{\left(\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)}{\sqrt{\alpha_{t}}}\right)}_{\text{ predicted }\hat{{\bm{z}}}_{0}^{(t)}}+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)}_{\text{direction pointing to }{\bm{z}}_{t}}+\underbrace{\sigma_{t}{\bm{\epsilon}}_{t}}_{\text{random noise}},$ (2)
@@ 4 Methodology
We first analyze the relation between optimization-based and reconstruction-based 3D scene editing methods in detail. Then, we elaborate on the proposed progressive Trajectory-Anchored Scheme (TAS) and the design of View-Consistent Attention Control (VCAC). The overall framework is illustrated in Figure 1.
4.1 Analysis on Optimization-based Editing
@@
$\hat{{\bm{z}}}_{0}^{(t)}\coloneq{\bm{f}}_{\phi}({\bm{z}}_{\pi},{\bm{\epsilon}},t,y)$ $=\frac{({\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y))}{\sqrt{\alpha_{t}}}$ (5)
@@
$\hat{{\bm{z}}}_{0}^{(t)}$ $=\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}}{\sqrt{\alpha_{t}}}=\frac{\sqrt{\alpha_{t}}\hat{{\bm{z}}}_{0}^{(t+1)}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{t+1}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}}{\sqrt{\alpha_{t}}}$ (6)
@@
We observe that the dynamics described by Eq. 6 closely resembles the optimization of SDS Eq. 5. With the assumption that SDS following an annealing timestep schedule, the SDS optimization is equivalent to an iterative reconstruction procedure: 1) sample next-step pseudo-ground-truth using noisy latents produced from previous rendered images with DDCM; 2) optimize a reconstruction loss to match the view with pseudo-ground-truth. This finding allows us to consider the design of 3D editing methods from a more general iterative reconstruction perspective. We expand this formulation to include variants of SDS, summarizing their pseudo-ground-truth parameterizations in Table LABEL:table:f-variants of the Appendix.
4.2 Trajectory-Anchored Scheme for Progressive 3D Gaussian Editing
@@
Our previous analysis in Sec. 4.1 bridges the gap between optimization-based and reconstruction-based 3D Gaussian editing methods, demonstrating that optimization-based approaches are special cases of reconstruction-based methods. Without loss of generality, we consider the design of 3D editing methods from a broader reconstruction-based perspective. The key questions for designing improved 3D Gaussian editing methods become: (1) which appropriate reconstruction pseudo-ground truths to use, and (2) how to schedule the reconstruction process of 3D Gaussians in a progressive manner for 3D editing.
@@
With this in mind, we propose a trajectory-anchored progressive 3D Gaussian editing scheme, as shown in Figure 1. This approach more tightly couples the image editing process with the rendering process, enabling 3D rendering to actively rectify the inconsistencies in the progressive 2D multi-view edits. Specifically, we use images generated by a modified DDCM editing trajectory, similar to InfEdit [34], as suitable pseudo-ground-truths for 3D Gaussian editing. The detailed sampling process for generating pseudo-ground-truths is described in Algorithm 1 in the Appendix. These pseudo-ground-truths enable a smooth transition from the source image to the target image (as illustrated in Figure 2), making them ideal for the progressive 2D editing and 3D updating scheme. Progressive edits on 2D views can be promptly applied to 3D Gaussians. Conversely, minor view inconsistencies arising from 2D editing can be promptly rectified through 3D constrained rendering.
@@
4.3 View-Consistent Attention Control

### v2

# TrAME: Trajectory-Anchored Multi-View Editing for Text-Guided 3D Gaussian Manipulation

@@ Abstract
Despite significant strides in the field of 3D scene editing, current methods encounter substantial challenge, particularly in preserving 3D consistency in multi-view editing process. To tackle this challenge, we propose a progressive 3D editing strategy that ensures multi-view consistency via a Trajectory-Anchored Scheme (TAS) with a dual-branch editing mechanism. Specifically, TAS facilitates a tightly coupled iterative process between 2D view editing and 3D updating, preventing error accumulation yielded from text-to-image process. Additionally, we explore the relationship between optimization-based methods and reconstruction-based methods, offering a unified perspective for selecting superior design choice, supporting the rationale behind the designed TAS. We further present a tuning-free View-Consistent Attention Control (VCAC) module that leverages cross-view semantic and geometric reference from the source branch to yield aligned views from the target branch during the editing of 2D views. To validate the effectiveness of our method, we analyze 2D examples to demonstrate the improved consistency with the VCAC module. Further extensive quantitative and qualitative results in text-guided 3D scene editing indicate that our method achieves superior editing quality compared to state-of-the-art methods. We will make the complete codebase publicly available following the conclusion of the review process.
Index Terms:
Diffusion Models, 3D Scene Editing, 3D Gaussian Splatting, Attention Mechanism
I Introduction
@@
3D technologies are pivotal in various fields, such as virtual reality (VR), the film and gaming industries, and 3D design [1, 2, 3, 4, 4, 5]. Recent advances in Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) have pushed the boundary of 3D representation, enabling versatile 3D applications [6, 7, 8, 9, 10, 11, 12, 13, 14]. Leveraging the powerful generative capabilities of text-to-image diffusion models [15, 16], text-guided 3D editing now allows for intricate adjustments in shape, style, texture, and lighting, marking a significant advancement in the flexibility of 3D scene manipulation [17, 18, 2, 3, 19, 20, 21, 22, 23].
@@
One key challenge in 3D editing is maintaining multi-view consistency when applied to real scenes [24, 25], which is essential for preventing visual artifacts and inconsistent appearances when observed from different view angles. Current studies of 3D editing can be divided into two categories, namely optimization-based and reconstruction-based methods. The optimization-based 3D editing methods [26, 27, 28, 29, 18] usually adopt the Score Distillation Sampling (SDS) loss [30] or its modified versions [31, 32]. They often suffer from sub-optimal editing quality such as over-saturation, over-smoothing, and a lack of diversity due to the misalignment between the randomly sampled timestep $t$ in the SDS optimization process and the corresponding timestep $t$ in the diffusion sampling process [33, 1]. The second category of methods, reconstruction-based 3D editing methods [17, 34, 25, 24, 19], employ readily available 2D editing diffusion models to edit 3D scene. However, 2D diffusion models face challenges in achieving multi-view consistency due to their intrinsic limitation of independently processing each view.
@@ II Related Work
Regarding optimization-based 3D editing, significant advancements have been made, building upon the foundational SDS loss introduced by DreamFusion [30]. This has catalyzed the development of various innovative methods [26, 27, 28, 29, 18, 32] aimed at refining 3D model editing via optimizing the SDS loss. Notably, Vox-E [27] and DreamEditor [29] have harnessed explicit 3D representations, including voxels and meshes. These approaches utilize cross-attention mechanisms to facilitate precise manipulations within specific regions of 3D models. Furthermore, RePaint-NeRF [38] has advanced the application of SDS in 3D editing by integrating a semantic mask to guide and constrain modifications within the background elements. In a similar vein, ED-NeRF [32] has introduced an enhanced loss function specifically designed for 3D editing tasks. It extends the Delta Denoising Score (DDS) loss [31] into the three-dimensional space, offering a more refined approach to editing. Posterior distillation sampling [26] aligns the identities of sampled targets with their corresponding sources by matching the stochastic latent variables obtained through DDPM inversion. Despite these advancements, a common challenge still persists among methods that utilize randomly sampled timesteps $t$ within the SDS framework: This approach tends to lead to deviations from the intended diffusion sampling trajectories, which can result in compromised quality of the 3D edits.
For reconstruction-based 3D editing, InstructN2N [17] alongside subsequent studies [34, 25, 24, 19, 39], have leveraged advancements in 2D diffusion editing techniques, notably InstructP2P [15] and ControlNet [40], for enhancing scene updates through Iterative Dataset Update [17]. GaussianEditor [34] and GaussCtrl [25] introduce methods for precise control in partial 3D editing. GaussianEditor utilizes semantic tracing to segment the 3D Gaussians to be edited. GaussCtrl proposes depth-guided editing and attention-based latent alignment. Despite their innovations, these methods often neglect inter-view correlations, resulting in a gap in maintaining semantic and structural consistency across multiple views, ultimately leading to error accumulation. ViCA-NeRF [41] adopts a depth-guided blending strategy to enforce 3D consistency across edited multi-views. However, this approach is prone to accumulating errors from depth estimation, frequently resulting in blurry and corrupted edits. Our work bridges this gap by integrating Trajectory-Anchored Scheme to minimize error accumulation and introducing attention manipulation mechanisms to enforce structural and semantic consistency across views in 3D editing through our VCAC module.
III Preliminaries
@@
III-A Score Distillation Sampling
@@
III-B Denoising Diffusion Consistent Model
@@
${\bm{z}}_{t-1}=$ $\sqrt{\alpha_{t-1}}\underbrace{\left(\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)}{\sqrt{\alpha_{t}}}\right)}_{\text{ predicted }\hat{{\bm{z}}}_{0}^{(t)}}$ (2)
$+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)}_{\text{direction pointing to }{\bm{z}}_{t}}+\underbrace{\sigma_{t}{\bm{\epsilon}}_{t}}_{\text{random noise}},$
@@ IV Methodology
We first analyze the relation between optimization-based and reconstruction-based 3D scene editing methods in detail. Then, we elaborate on the proposed progressive Trajectory-Anchored Scheme (TAS) and the design of View-Consistent Attention Control (VCAC). The overall framework is illustrated in Fig. 1.
IV-A Analysis on Optimization-based Editing
@@
$\hat{{\bm{z}}}_{0}^{(t)}$ :-$\coloneq{\bm{f}}_{\phi}({\bm{z}}_{\pi},{\bm{\epsilon}},t,y):$ (5)
$=\frac{({\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y))}{\sqrt{\alpha_{t}}}$
@@
$\hat{{\bm{z}}}_{0}^{(t)}$ $=\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}}{\sqrt{\alpha_{t}}}$ (6)
$=\frac{\sqrt{\alpha_{t}}\hat{{\bm{z}}}_{0}^{(t+1)}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{t+1}-\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{\phi}}{\sqrt{\alpha_{t}}}$
@@
We observe that the dynamics described by Eq. 6 closely resembles the optimization of SDS Eq. 5. With the assumption that SDS following an annealing timestep schedule, the SDS optimization is equivalent to an iterative reconstruction procedure: 1) sample next-step pseudo-ground-truth using noisy latents produced from previous rendered images with DDCM; 2) optimize a reconstruction loss to match the view with pseudo-ground-truth. This finding allows us to consider the design of 3D editing methods from a more general iterative reconstruction perspective. We expand this formulation to include variants of SDS, summarizing their pseudo-ground-truth parameterizations in Table I.
TABLE I: Denote $\gamma_{t}=\sqrt{1-\alpha_{t}}/\sqrt{\alpha_{t}}$, we compare different pseudo-ground-truths of SDS and its variants.
Methods Pseudo-ground-truths ${\bm{f}}$
SDS / SJC [30, 42] ${\bm{z}}_{\pi}+\gamma_{t}\left({\bm{\epsilon}}-{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)\right)$
VSD [46] ${\bm{z}}_{\pi}+\gamma_{t}\left({\bm{\epsilon}}_{\varphi}({\bm{z}}_{t},t,y)-{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)\right)$
DDS [31] ${\bm{z}}_{\pi}+\gamma_{t}\left({\bm{\epsilon}}_{\phi}({\bm{z}}_{t}^{\prime},t,y^{\prime})-{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)\right)$
ISM [47] ${\bm{z}}_{\pi}+\gamma_{t}\left({\bm{\epsilon}}_{\phi}({\bm{z}}_{s},s,\varnothing)-{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y)\right)$
NFSD [48] ${\bm{z}}_{\pi}+\gamma_{t}\left({\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,y_{neg})-{\bm{\epsilon}}_{\phi}({\bm{z}}_{t},t,\varnothing)-s{\bm{\delta}}_{\phi}^{C}\right)$
IV-B Trajectory-Anchored Scheme for Progressive 3D Gaussian Editing
@@
Our previous analysis in Sec. IV-A bridges the gap between optimization-based and reconstruction-based 3D Gaussian editing methods, demonstrating that optimization-based approaches are special cases of reconstruction-based methods. Without loss of generality, we consider the design of 3D editing methods from a broader reconstruction-based perspective. The key questions for designing improved 3D Gaussian editing methods become: (1) which appropriate reconstruction pseudo-ground truths to use, and (2) how to schedule the reconstruction process of 3D Gaussians in a progressive manner for 3D editing.
@@
With this in mind, we propose a trajectory-anchored progressive 3D Gaussian editing scheme, as shown in Fig. 1. This approach more tightly couples the image editing process with the rendering process, enabling 3D rendering to actively rectify the inconsistencies in the progressive 2D multi-view edits. Specifically, we use images generated by a modified DDCM editing trajectory, similar to InfEdit [35], as suitable pseudo-ground-truths for 3D Gaussian editing. Consider the DDCM pseudo-inversion process in InfEdit [35]:
${\bm{z}}^{\text{tgt}}_{0}\coloneq\frac{{\bm{z}}^{\text{tgt}}_{t}-\sqrt{1-\alpha_{t}}({\bm{\epsilon}}^{\text{tgt}}_{\phi}-{\bm{\epsilon}}^{\text{src}}_{\phi}+\epsilon)}{\sqrt{\alpha_{t}}}$ (7)
where ${\bm{\epsilon}}^{\text{src}}_{\phi}$ and ${\bm{\epsilon}}^{\text{tgt}}_{\phi}$ are the noise prediction from the source branch and target branch, respectively. Let
$&\hat{{\bm{z}}}^{\text{src}}_{0}\coloneq\frac{{\bm{z}}^{\text{src}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}^{\text{src}}_{\phi}}{\sqrt{\alpha_{t_{n}}}}\\ &\hat{{\bm{z}}}^{\text{tgt}}_{0}\coloneq\frac{{\bm{z}}^{\text{tgt}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}^{\text{tgt}}_{\phi}}{\sqrt{\alpha_{t_{n}}}}\end{aligned}\;,$ (8)
the pseudo-ground-truth ${\bm{z}}^{\text{tgt}}_{0}$ can be reparameterized as
${\bm{z}}^{\text{tgt}}_{0}=\hat{{\bm{z}}}^{\text{tgt}}_{0}+\kappa\left({\bm{z}}^{\text{src}}_{0}-\hat{{\bm{z}}}^{\text{src}}_{0}\right)\end{aligned}\;.$ (9)
Here, $\kappa$ denotes a DDCM adjustment coefficient. When $\kappa=1$, the pseudo-ground-truth trajectory coincides with the editing trajectory of InfEdit. We observe that the second term in Eq. 9 injects information from the original image into $\hat{{\bm{z}}}^{\text{tgt}}_{0}$ during the editing process. Consequently, a large $\kappa$ may lead to an “overshooting” effect, resulting in over-saturated color and over-sharpened edges, analogous to image sharpening. In our experiments, we evaluate the impact of various values of $\kappa$ on 2D view editing and 3D scene editing and select the optimal value for our method.
The detailed sampling process for generating pseudo-ground-truths is described in Algorithm 1. These pseudo-ground-truths enable a smooth transition from the source image to the target image (as illustrated in Fig. 2), making them ideal for the progressive 2D editing and 3D updating scheme. Progressive edits on 2D views can be promptly applied to 3D Gaussians. Conversely, minor view inconsistencies arising from 2D editing can be promptly rectified through 3D constrained rendering.
@@
Algorithm 1 Trajectory-Anchored Scheme
Input: $\theta^{\text{src}}$, $\left\{t_{n}\right\}_{n=1}^{N}$, $y^{\text{src}}$, $y^{\text{tgt}}$, $\eta$, $K$, $\kappa$
initialize $\theta^{(t_{n})}\leftarrow\theta^{\text{src}}$.
for $n=1$ to $N$ do
     Sample ${\bm{\epsilon}}\sim\mathcal{N}(\bm{0},{\bm{I}})$
     ${\bm{z}}_{\pi}\leftarrow{\bm{g}}(\theta^{(t_{n})})$
     ${\bm{z}}^{\text{tgt}}_{t_{n}}\leftarrow\sqrt{\alpha_{t_{n}}}{\bm{z}}_{\pi}+\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}$
     ${\bm{z}}^{\text{src}}_{t_{n}}\leftarrow\sqrt{\alpha_{t_{n}}}{\bm{z}}^{\text{src}}_{0}+\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}$
     $\hat{{\bm{z}}}^{\text{src}}_{0}\leftarrow\left({\bm{z}}^{\text{src}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}^{\text{src}}_{\phi}\right)/\sqrt{\alpha_{t_{n}}}$
     $\hat{{\bm{z}}}^{\text{tgt}}_{0}\leftarrow\left({\bm{z}}^{\text{tgt}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}}{\bm{\epsilon}}^{\text{tgt}}_{\phi}\right)/\sqrt{\alpha_{t_{n}}}$
     ${\bm{z}}^{\text{tgt}}_{0}\leftarrow\hat{{\bm{z}}}^{\text{tgt}}_{0}+\kappa\left({\bm{z}}^{\text{src}}_{0}-\hat{{\bm{z}}}^{\text{src}}_{0}\right)$
     for $k=1$ to $K$ do
         ${\bm{z}}_{\pi}\leftarrow{\bm{g}}(\theta^{(t_{n})})$
         $\theta^{(t_{n})}\leftarrow\theta^{(t_{n})}-\eta\nabla_{\theta}\mathcal{L}({\bm{z}}_{\pi},{\bm{z}}^{\text{tgt}}_{0})$
     end for
     $\theta^{(t_{n-1})}\leftarrow\theta^{(t_{n})}$
end for
return $\theta^{(t_{1})}$
IV-C View-Consistent Attention Control